OpenStudy (anonymous):

HELP PLEASE , ILL MEDAL Determine the number of solutions for the system of equations. 2x + 8y = 6 -5x - 20y = -15 A. No Solutions B. One Solution C. Exactly 2 solutions D. Infinite number of solutions

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@coolman500

OpenStudy (welshfella):

Hint :- multiply the first equation by 5 and the second equation by 2 what do you notice?

OpenStudy (welshfella):

- rather multiply second by -2

OpenStudy (welshfella):

2x * 5 -8y*5 = 6*5 10x - 40y = 30

OpenStudy (welshfella):

sorry that's 10x + 40y = 30 (its +8y not -8y)

OpenStudy (welshfella):

now multiply the second equation by -2

OpenStudy (welshfella):

can you do that?

OpenStudy (welshfella):

-5x*-2 - 20y*-2 = -15*-2 can you continue?

OpenStudy (welshfella):

remember - * - = +

OpenStudy (welshfella):

r u there?